# C2P: Featuring Large Language Models with Causal Reasoning

### v2

@@ 4.3 Evaluation of the C2P on natural stories
GPTs
GPT-4.5 2 2 3 3
GPT-4 2 2 3 3
GPT-4o 3 2 3 3
Featured GPT
C2P on GPT-4 Turbo 5 0 5 0
Few-shot learned GPT-4 Turbo 4 0 5 1
Additionally, Table 4 provides an example of such natural stories and demonstrates the result of applying a few-shot learned GP4-Turbo with C2P to it. This showcases how GPT models (including versions 3.5, 4, 4 Turbo, and 4o) fail to respond correctly. It is important to note that the few-shot learning process of GPT-4 Turbo was performed using only six examples from the CORR2CAUSE dataset, without including any natural story examples. Note that there are no natural stories employed in the few-shot learning process.
Table 4: Example of a natural story and the result of few-shot learned GPT-4-Turbo with C2P
Example of true story Results of few-shot learned GPT-4 Turbo with C2P
Premise: eating junk food (A), obesity (C), and watching television (B) have relations with each other. There is a correlation between eating junk food and obesity, and between watching television and obesity. However, eating junk food and watching television are independent from each other. Hypothesis: Eating junk food directly affects obesity Step 1: Number of random variable:3 Names of random variable: A, B, C
Step 2: All of Statistical Relations:{ Dependencies: [[A, C], [B, C]], Unconditional Independencies: [[A, B]], Conditional Independencies: []}
Step 3: A: {A: 0, B: 1, C: 1}, B: {A: 1, B: 0, C: 1}, C: {A: 1, B: 1, C: 0}
Step 4: A: {A: 0, B: 0, C: 1}, B: {A: 0, B: 0, C: 1}, C: {A: 1, B: 1, C: 0}
Step 5: A: {A: 0, B: 0, C: 1}, B: {A: 0, B: 0, C: 1}, C: {A: 1, B: 1, C: 0}
@@
Step 9: Checking matrix[A][D] = 1 and matrix[D][A] = 0. According to rule 2, this suggests A is a direct cause of D, or D is a direct effect of A.
@@
4.4 Evaluation of the C2P framework on the (co)evolution of supermassive black holes and their host galaxies
@@
In astrophysics, experiments are impossible. Thus, causal reasoning must rely exclusively on observational data. Similar to the study in (Pasquato et al., 2023), we (co)evolute supermassive black holes (SMBHs) and their host galaxies. We implement C2P on GPT-4 Turbo to a comprehensive catalog of galaxy properties to obtain a PDAG and then reason the causal questions.
Fig. 3 shows the results of each sub-tasks of implementing C2P
@@
Figure 3: Prompts (Q) and results (A) of sub-tasks application of the C2P framework to real-world complex scenarios and steps of sub-task 3 for the given premise.
@@
According to the extracted PDAG, many questions such as the causal given can be answered which is too complicated for the existing LLMs to answer. It is important to note that by implementing C2P, not only many causal questions can be answered with this approach, but it is easy to show which causal questions can not be answered based on the given premise and which more information has to be given to be able to causally reason. More importantly, it is worth noting that this is not the limitation of the C2P framework, any rational agent needs such information to answer such causal questions.
5 Discussion, Limitations and Future Work
@@
Causal reasoning and identifiability: Having a PDAG of an SCM of a causal mechanism is the most effective way to achieve this with observational data. Based on this, two main questions arise: Given a PDAG, under what conditions can we make causal reasoning? This involves determining the necessary assumptions and data required to address a causal question—this issue is known as identification. It is generally possible to answer some cause-effect questions based on the PDAG, represented by a causal graph, and lower-level data. These questions are mainly the ones that are related to the part where we have extracted a collider, or if the question is just about the existence of a cause-effect relation; in some cases, these questions can be answered as well. However, it is almost impossible to answer all the causal questions based on the given information, as discussed in (Bareinboim et al., 2022; Pearl & Bareinboim, 2022). As a result, the second question is, what else is needed to answer all the causal questions? In such cases, interventional data is necessary to explicitly extract the direction of causal relations in a mechanism. To do so, while a naive baseline approach would require $O(n^{2})$ interventions, various methods have been proposed, such as those in (Kocaoglu et al., 2017; Choo & Shiragur, 2023; von Kügelgen et al., 2024; Squires & Uhler, 2023), for cases in different situations. However, one clear thing is that it is not possible to reason all the causal questions. For instance,
@@
Causal reasoning is a broad and complex task in AI and LLMs. While current machine learning methods struggle with extracting causal structures and subsequently reasoning causally, this problem is even more intricate in the context of LLMs. The primary reason for this complexity is the inherent structure of LLMs and their reliance on the concept of attention. To enable LLMs to reason causally, the first step is to identify the causal question. Recently, a comprehensive database was compiled, and an AI model was trained to identify causal questions (Ceraolo et al., 2024). This model can be used prior to applying our framework to determine whether a causal reasoning process has to be activated or not. Additionally, more comprehensive examples and scenarios need to be generated to aid in the learning process of an LLM. Due to the vast variety of causal scenarios and the token limitations of current models, our approach must be integrated into the learning or fine-tuning process of an LLM. This ensures that the model can effectively learn and apply causal reasoning within its capabilities.
Scaling Instruction-Finetuned Language Models: LLMs primarily function through pattern recognition without the capability to recognize causal relationships, as indicated by prior research . This limitation highlights the importance of implementing these models with causal analysis techniques to enhance their causal reasoning ability. Previous attempts to boost LLMs’ causal reasoning involved using external tools or manually crafted prompts, which, although insightful, are impractical for easy integration. Our study introduced an algorithm that leverages LLMs’ inherent abilities to foster causal reasoning autonomously. Our findings demonstrate a marked increase in the accuracy of LLMs’ causal understanding, even in scenarios involving complex causal interactions. Additionally, incorporating a chain of thought into LLMs’ training has been proven to enhance performance across various tasks (Chung et al., 2024). Based on the extracted results on the benchmark and provided datasets in our study, and the results of the few-shot learning process, we strongly hypothesize that our approach could similarly benefit the training of language models, particularly in boosting their causal reasoning, even in smaller models that would lead to a revelation in LLMs. This enhancement could be realized using a carefully selected set of examples or synthetic samples created by LLMs, following the guidelines outlined in our algorithm.
Future works As demonstrated in the results section, C2P has significantly enhanced the reasoning capabilities of LLMs. We have shown that extracting the adjacency matrix of the variables in a given premise can function as the DAG in Pearl’s works, leading to improved reasoning ability. Our experimental results of few-shot learned LLMs indicate a significant increase in accuracy for tasks where users inquire about direct cause-and-effect relationships or common cause questions. However, the models still struggle with indirect cause-and-effect questions. Pearl demonstrated that the concepts of a DAG and d-separation complement each other in causal reasoning tasks. To address the low accuracy in answering indirect causal reasoning questions, more sub-tasks can be introduced to LLMs to perform d-separation in a more structured manner. As mentioned, due to the token limitation of LLMs, we were able to use only six or even fewer examples in the few-shot process of GPT-4-Turbo. As a result, our proposed framework can be seamlessly integrated into the fine-tuning or learning process of an LLM by introducing more extensive examples with more complex scenarios. Given the accuracy achieved in few-shot learned LLMs, the fine-tuned or trained LLMs with our approach can achieve highly accurate results, effectively addressing the causal reasoning bottleneck of LLMs.
@@ Appendix A Appendix
A.1 Prompts for C2P frameworks
The proposed prompts for step-by-step proceeding with C2P are as follows:
@@ A.1 Prompts for C2P frameworks
    1. (i)
      Please give me the number of random variables in the given premise
    2. (ii)
      Write the names of all random variables.
@@
    • If 2 random variables, for instance, $R_{1}$ and $R_{2}$, are independent, write it in this form: "$R_{1}$ is independent of $R_{2}$". If there exist 2 random variables, for instance, $R_{1}$ and $R_{2}$, are conditionally independent given a third random variable, for instance, $R_{3}$, write it in this form: "$R_{1}$ and $R_{2}$ are independent given $R_{3}$". If two random variables, for instance, $R_{1}$ and $R_{2}$, are specially mentioned to have cause and effect relation, write it in this form: "$R_{1}$ is the cause of $R_{2}$".
@@
    • In this phase, each random variable is treated as a node within a fully connected undirected graph. Then, for each pair, for instance, $R_{1}$ and $R_{2}$, presented in the form: "$R_{1}$ is the cause of $R_{2}$", set the element in ["$R_{2}$", "$R_{1}$"] in the adjacency matrix to 0.
@@
    1. (i)
      Update the adjacency matrix based on the specified unconditional independencies between random variables. Each pair of variables that is declared independent should have their corresponding value set to zero in the adjacency matrix. The initial adjacency matrix and the list of independencies are provided below. Please ensure all independencies are correctly reflected in the updated matrix. Instructions: - For each pair of variables listed as independent, set their corresponding entries in the adjacency matrix to 0.
    2. (ii)
      Update the adjacency matrix based on the specified conditional independencies between random variables. Each pair of variables that is declared independent should have their corresponding value set to zero in the adjacency matrix. The initial adjacency matrix and the list of independencies are provided below. Please ensure all independencies are correctly reflected in the updated matrix. Instructions: - For each pair of variables listed as independent given other variable(s), set their corresponding entries in the adjacency matrix to 0.
    3. (iii)
      Task: Given an initial adjacency matrix, follow these steps: Step 1: Identify all rows (key values) in the matrix where there are two or more than two columns with the value "1" in them. For each identified row, find all pairs of different columns where the values are "1".Ensure to exclude rows that do not contain any pairs from the results. Step 2: Display these pairs, "All Pairs", where each row name is key, and the value is a list of column names that are identified in Step 1.
    4. (iv)
      Given the "All Pairs" and the list of independencies, follow these instructions step by step: Instruction: For each key in "All Pairs", delete all the pairs that are not mentioned as independent in the "independencies" list and return other with all their values. The "All Pairs" contains pairs of elements associated with each key. The goal is to update this by removing pairs that are not mentioned as independent. The list of independencies provides information about which pairs are independent of each other.
    5. (v)
      Given the initial adjacency matrix represented and the "All Pairs" list, for each key-value pair ("$R$") in "All Pairs", modify the initial adjacency matrix as follows: 1- Set the value in the "$C_{1}$" row and "$R$" column to 0: ("$C_{1}$", "$R$") = 0. 2- Set the value in the "$C_{2}$" row and "$R$" column to 0: ("$C_{2}$", "$R$") = 0.
@@
    • To extract and understand causal relations in the adjacency matrix: For each specified variable "$R$" and "$C$", for instance, that are listed in the adjacency matrix: - If matrix entry at ["$R$", "$C$"] = 1 and ["$C$", "$R$"] = 1, then the causal direction between "$R$" and "$C$" is undetermined. - If matrix entry at ["$R$", "$C$"] = 1 and ["$C$", "$R$"] = 0, then "$R$" is a direct cause of "$C$" or "$C$" is a direct effect of "$R$". - If matrix entry at ["$R$", "$C$"] = 0 and ["$C$", "$R$"] = 1, then "$C$" is a direct cause of "$R$" or "$R$" is a direct effect of "$C$". If two variables directly affect a third variable, the first two variables are common causes, and the third variable is a collider. Evaluate the hypothesis based on the given partially directed acyclic graph presented as an adjacency matrix with the given Instruction.
A.2 Prompts for few-shot learning of C2P
An example of given prompts for the few-shot learning process of C2P is as follows:

### v4

@@ 4.4 Evaluation of the $\text{C}^{2}\text{P}$ on natural stories
LLMs without $\text{C}^{2}\text{P}$
GPT-3.5 5 8 7 10
GPT-4 Turbo 6 7 8 9
GPT-4 7 8 7 8
GPT-4O 8 6 9 7
LLaMA 3-70B 9 9 6 6
LLaMA 3.1-70B 9 10 5 6
Featured LLMs with $\text{C}^{2}\text{P}$
$\text{C}^{2}\text{P}$ on LLaMA 3.1-70B 11 3 12 4
$\text{C}^{2}\text{P}$ Few-shot learned LLaMA 3.1-70B 10 3 12 5
$\text{C}^{2}\text{P}$ on GPT-4 Turbo 13 1 14 2
$\text{C}^{2}\text{P}$ Few-shot learned GPT-4 Turbo 11 3 12 4
Additionally, Tab. 5 provides an example of such natural stories and demonstrates the result of applying few-shot learned GP4-Turbo and LLaMA 3.1 with $\text{C}^{2}\text{P}$ to them. This showcases how GPT models (including versions 3.5, 4, 4 Turbo, and 4O) fail to respond correctly. It is important to note that the few-shot learning process of GPT-4 Turbo and LLaMA 3.1 was performed using only ten examples from the CORR2CAUSE dataset, without including any natural story examples.
Table 5: Example of a natural story and the result of few-shot learned GPT-4-Turbo with $\text{C}^{2}\text{P}$
Example of true story Results of few-shot learned GPT-4 Turbo with $\text{C}^{2}\text{P}$
Premise: eating junk food (A), obesity (C), and watching television (B) have relations with each other. There is a correlation between eating junk food and obesity, and between watching television and obesity. However, eating junk food and watching television are independent from each other. Hypothesis: Eating junk food directly affects obesity Step 1: Number of random variable:3
Names of random variable: A, B, C
Step 2: All of Statistical Relations:{
Dependencies: [[A, C], [B, C]],
Unconditional Independencies: [[A, B]],
Conditional Independencies: []}
Step 3: A: {A: 0, B: 1, C: 1},
B: {A: 1, B: 0, C: 1},
C: {A: 1, B: 1, C: 0}
Step 4: A: {A: 0, B: 0, C: 1},
B: {A: 0, B: 0, C: 1},
C: {A: 1, B: 1, C: 0}
Step 5: A: {A: 0, B: 0, C: 1},
B: {A: 0, B: 0, C: 1},
C: {A: 1, B: 1, C: 0}
@@
Step 9: Checking matrix[A][C] = 1 and matrix[C][A] = 0. According to rule 2, this suggests A is a direct cause of C, or C is a direct effect of A.
@@
4.5 Evaluation of the $\text{C}^{2}\text{P}$ framework on the (co)evolution of supermassive black holes and their host galaxies
@@
In astrophysics, interventional experiments are impossible. Thus, causal reasoning must rely exclusively on observational data. Pasquato et al. (2023) computationally studied the coevolution of supermassive black holes based on numerical data, and a causal graph for the underlying mechanism is extracted. Then, the causal hypothesis, "central density affects black hole mass", is answered. Similar to their study, we (co)evolute supermassive black holes (SMBHs) and their host galaxies based on the verbalized information in (Pasquato et al., 2023). We GPT-4 Turbo with $\text{C}^{2}\text{P}$ to process the verbalized information provided in subsection 4.1, obtain a PDAG, and then reason about the causal questions.
Fig. 4 illustrates the results of the subtasks of $\text{C}^{2}\text{P}$ on SMBHs data presented.
@@
Figure 4: Prompts (Q) and results (A) of subtasks application of the $\text{C}^{2}\text{P}$ framework to real-world complex scenarios and steps of subtask 3 for the given premise.
@@
The first four subtasks aim to extract the PDAG from the given data. Then, the hypothesis that central density affects black hole mass is evaluated in subtask 5 and answered. According to the extracted PDAG, many other reasoning questions questions can be answered that are too complicated for the existing LLMs to answer. It is important to note that by implementing $\text{C}^{2}\text{P}$, not only many causal questions can be answered with this approach, but it is easy to show which causal questions can not be answered based on the given premise and which more information has to be given to be able to causally reason. The discussion on which questions can and cannot be answered based on PDAG is presented in detail in studies such as (Hernán & Robins, 2006; Hauser & Bühlmann, 2012; Perkovic, 2020). Importantly, the inability to answer certain causal reasoning questions based on the PDAG, or its corresponding adjacency matrix, does not reflect a limitation of the $\text{C}^{2}\text{P}$ framework. Rather, any rational agent needs additional information to effectively address all causal inquiries.
5 Discussion, Challenges and Future Work
Practical insights on simulations and results of $\text{C}^{2}\text{P}$: Each part of our study has a specific aim. The goal of the simulations, where we applied the 9 steps of $\text{C}^{2}\text{P}$ step by step, is to show what the expected output would be and how it leads to causal reasoning. However, our investigations demonstrate sensitivity to the prompts. In few-shot learning, where examples of expected results are provided for each prompt, the sensitivity to prompts was not significant; instead, the results were primarily dependent on the number of examples. Interestingly, while the prompt is lengthy —due to the fact that all steps are given to LLMs as a single prompt —- the results remained highly accurate with as few as ten examples, as reported in Tab.s LABEL:table4 and 5. By comparing the prompts in Appendices A.3 and A.4, it is clear that the prompts in the few-shot learning section do not need to be as detailed and are more abstract than those used in the step-by-step application of $\text{C}^{2}\text{P}$. Additionally, it is important to note that the results for natural stories using few-shot learned $\text{C}^{2}\text{P}$ were highly accurate (see Table 5), even though the examples in the few-shot learning process were synthetic “cause-effect" examples from the CORR2CAUSE dataset rather than real story examples.
The comparison of results between $\text{C}^{2}\text{P}$ and CoT provides new insights into how LLMs aim to reason and why $\text{C}^{2}\text{P}$ and similar frameworks can guide a model to "think" step by step. As demonstrated in the examples provided in the code repository—under "Sample Responses of LlaMa 3.1 with CoT" for LlaMa 3.1 and "Sample Responses of GPT-4 Turbo with CoT" for GPT-4 Turbo— both models attempt to break tasks into multiple subtasks when they are asked to think step by step (sometimes with more than nine consecutive subtasks). However, they still lack rationality in how they structure these subtasks, as it is discussed in Wei et al. (2022a). In other words, they struggle to design the necessary subtasks for reasoning, a task that $\text{C}^{2}\text{P}$ successfully performs to enhance reasoning. This challenge is also highlighted in Wei et al. (2022b), where CoT is introduced, and its logical inconsistencies and poor step alignment are discussed as key limitations of the approach.
@@ 5 Discussion, Challenges and Future Work
Causal reasoning and identifiability: Extracting a PDAG using only observational data is a key step in learning the true underlying causal mechanism. Based on the extracted PDAG, two main questions arise: Given a PDAG, under what conditions can we make causal reasoning? This involves determining the necessary assumptions and data required to address a causal question. This issue is known as identifiability. It is generally possible to answer some cause-effect questions based on the PDAG, represented by a causal graph, and lower-level data. These questions are mainly the ones that are related to the part where we have extracted a collider, or if the question is just about the existence of a cause-effect relation; in some cases, these questions can be answered as well. However, it is almost impossible to answer all the causal questions based on the given information, as discussed in (Bareinboim et al., 2022; Pearl & Bareinboim, 2022). As a result, the second question is, what else is needed to answer all the causal questions? In such cases, interventional data is necessary to explicitly extract all the directions of causal relations in a mechanism. To do so, while a naive baseline approach would require $O(n^{2})$ interventions, various methods have been proposed, such as those in (Kocaoglu et al., 2017; Choo & Shiragur, 2023; von Kügelgen et al., 2024; Squires & Uhler, 2023), for cases in different situations. However, one clear thing is that it is not possible to reason all the causal questions. For instance,
@@
Next Steps for Reasoning in LLMs: The human brain possesses an intrinsic drive to understand causality. Whether driven by curiosity or the pursuit of specific goals, we continuously seek to understand why events occur and how they are interconnected. Causal reasoning is a broad and complex task in AI and LLMs. While current machine learning methods find it difficult to extract causal structures and subsequently reason causally, this problem can be even more intricate within the context of LLMs. The primary reason for this complexity is the distinction between association and cause-effect relationships. The inherent structure of LLMs relies on the attention mechanism presented in (Vaswani et al., 2017), which is akin to what Pearl refers to as "association" in the ladder of causality (Pearl & Mackenzie, 2018). As a result, advancing up the ladder of causality is essential for enabling true reasoning. To enable LLMs to address this fundamental quest for causality, they must be capable of sequentially performing two main tasks. As demonstrated in our study, $\text{C}^{2}\text{P}$ has significantly enhanced the reasoning capabilities of LLMs on the provided datasets, which can be the main task in the reasoning process. However, the first task is to be able to understand the causal questions, which involves formalizing the definition of causal questions and establishing a taxonomy for finer-grained classification. The initial effort on this is presented in Ceraolo et al. (2024). Then, $\text{C}^{2}\text{P}$ can be employed. However, the first subtask of this framework is to extract random variables, which requires the models to understand the definition of a random variable to ensure that the variables in the reasoning process are correctly identified. Failure in this process could undermine our proposed framework. Another important concept in this process is extracting the dependencies provided in the premise. In the datasets used for our experiments, these dependencies are explicitly stated in the given premise; however, in reality, they are often implicit. As a result, LLMs should be able to accurately extract these dependencies. Additionally, more comprehensive examples and scenarios need to be generated to aid in the learning process of an LLM. By overcoming these challenges, the integration of $\text{C}^{2}\text{P}$ with LLMs can provide these models with causal reasoning capabilities, similar to the transformative impact of "Chain-of-Thought" (Wei et al., 2022b), as highlighted by Chung et al. (2024). It is also important to note that while our experimental results with few-shot learned LLMs show a significant increase in accuracy for tasks involving direct cause-and-effect relationships or common cause questions, the models still struggle with indirect cause-and-effect questions. Moreover, the accuracy of our framework decreases as the number of random variables increases. As Pearl demonstrated that DAGs and d-separation are complementary in causal reasoning tasks, more structured subtasks can be introduced to improve performance in answering indirect causal questions and handling more complex scenarios with numerous variables. These subtasks would help guide LLMs in performing reasoning tasks more effectively.
The key benefit of few-shot learning is that it demonstrates that pre-trained models can perform causal reasoning without needing to be retrained from scratch or necessarily fine-tuned, making it more cost-effective and compute-efficient. This leads to reduced costs and less demanding infrastructure requirements. Therefore, a model fully trained or fine-tuned with $\text{C}^{2}\text{P}$ would probably perform even better than few-shot learned ones, as it can be trained or fine-tuned on thousands of examples instead of the ten examples used during few-shot learning. This highlights how integrating $\text{C}^{2}\text{P}$ during the training and fine-tuning of LLMs can revolutionize existing models. This is also discussed in Chung et al. (2024) as the “scaling law”. Additionally, Studies like Kaplan et al., 2020 showed that language models benefit from increased data and computational resources. More data during fine-tuning allows the model to capture better representations, leading to improved performance in various tasks. This is one of the core principles behind why larger datasets, would likely boost accuracy compared to just 10 examples. In Brown et al. (2020) with GPT-3, the authors highlight how performance increases as the model is trained on more examples. In few-shot learning, the model can generalize from a small set of examples, but fine-tuning with larger datasets usually leads to more substantial improvements. Devlin (2018) also demonstrates that fine-tuning with more task-specific examples improves the model’s performance. As a result, by addressing the mentioned challenges, employing $\text{C}^{2}\text{P}$ in the fine-tuning process is expected to improve the reasoning accuracy of the models.
@@ Appendix A Appendix
A.1 do-calculus rules
The three main rules of the do-calculus are as follows:
  1. 1.
  2. 2.
  3. 3.
A.2 Example of comparison in responding to causal queries
For the given example in Subsection 4.1, Tab. LABEL:tab:_wrong_and_right demonstrates that while GPT-4 arrives at the correct final answer, the reasoning process is flawed, providing further evidence of the random behavior of existing models in causal reasoning.
Table 6: An example of the correct final answer with GPT-4 Turbo while the reasoning process leading is incorrect and results of $\text{C}^{2}\text{P}$ Few-shot learned GPT-4 Turbo
GPT-4 Turbo with CoT $\text{C}^{2}\text{P}$ Few-shot learned GPT-4 Turbo
The hypothesis that A directly affects C could be possible. The premise states that A correlates with C, which means there is a relationship between them. However, correlation does not necessarily imply causation. It could be that A affects C, C affects A, or that both A and C are affected by a separate, unmentioned variable. However, since the premise does not provide any information contradicting the hypothesis, it is possible that A directly affects C. Final Answer: "Yes" According to the adjacency matrix, matrix["A"]["C"] = 1 and matrix["C"]["A"] = 0 indicates that A directly cause C. Therefore, the hypothesis that A causes C or C is effect of A is correct. Final Answer: "Yes"
A.3 Prompts for $\text{C}^{2}\text{P}$ frameworks
The proposed steps (prompts) for step-by-step proceeding with 5 main subtasks of $\text{C}^{2}\text{P}$ are as follows:
@@ A.3 Prompts for $\text{C}^{2}\text{P}$ frameworks
    • Step 1: Please give the number of random variables in the given premise and write the names of all random variables.
@@
    • Step 2: If 2 random variables, for instance, $R_{1}$ and $R_{2}$, are independent, write it in this form: $R_{1}$ is independent of $R_{2}$". If there exist 2 random variables, for instance, $R_{1}$ and $R_{2}$, are conditionally independent given a third random variable, for instance, $R_{3}$, write it in this form: $R_{1}$ and $R_{2}$ are independent given $R_{3}$". If two random variables, for instance, $R_{1}$ and $R_{2}$, are specially mentioned to have cause and effect relation, write it in this form: $R_{1}$ is the cause of $R_{2}$".
@@
    • Step 3: In this phase, each random variable is treated as a node within a fully connected undirected graph. Then, for each pair, for instance, $R_{1}$ and $R_{2}$, presented in the form: $R_{1}$ is the cause of $R_{2}$", set the element in $R_{2}$", $R_{1}$"] in the adjacency matrix to 0.
@@
    • Step 4: Update the adjacency matrix based on the specified unconditional independencies between random variables. Each pair of variables that is declared independent should have their corresponding value set to zero in the adjacency matrix. The initial adjacency matrix and the list of independencies are provided below. Please ensure all independencies are correctly reflected in the updated matrix. Instructions: - For each pair of variables listed as independent, set their corresponding entries in the adjacency matrix to 0.
    • Step 5: Update the adjacency matrix based on the specified conditional independencies between random variables. Each pair of variables that is declared independent should have their corresponding value set to zero in the adjacency matrix. The initial adjacency matrix and the list of independencies are provided below. Please ensure all independencies are correctly reflected in the updated matrix. Instructions: - For each pair of variables listed as independent given other variable(s), set their corresponding entries in the adjacency matrix to 0.
    • Step 6: Task: Given an initial adjacency matrix, follow these steps: Step 1: Identify all rows (key values) in the matrix where there are two or more than two columns with the value "1" in them. For each identified row, find all pairs of different columns where the values are "1".Ensure to exclude rows that do not contain any pairs from the results. Step 2: Display these pairs, "All Pairs", where each row name is key, and the value is a list of column names that are identified in Step 1.
    • Step 7: Given the "All Pairs" and the list of independencies, follow these instructions step by step: Instruction: For each key in "All Pairs", delete all the pairs that are not mentioned as independent in the "independencies" list and return other with all their values. The "All Pairs" contains pairs of elements associated with each key. The goal is to update this by removing pairs that are not mentioned as independent. The list of independencies provides information about which pairs are independent of each other.
    • Step 8: Given the initial adjacency matrix represented and the "All Pairs" list, for each key-value pair ("RR") in "All Pairs", modify the initial adjacency matrix as follows: 1- Set the value in the $C_{1}$" row and "RR" column to 0: ($C_{1}$", "RR") = 0. 2- Set the value in the $C_{2}$" row and "RR" column to 0: ($C_{2}$", "RR") = 0.
@@
    • Step 9: To extract and understand causal relations in the adjacency matrix: For each specified variable "RR" and "CC", for instance, that are listed in the adjacency matrix: - If matrix entry at [ "RR", "CC"] = 1 and ["CC", "RR"] = 1, then the causal direction between "RR" and "CC" is undetermined. - If matrix entry at ["RR", "CC"] = 1 and ["CC", "RR"] = 0, then "RR" is a direct cause of "CC" or "CC" is a direct effect of "RR". - If matrix entry at ["RR", "CC"] = 0 and ["CC", "RR"] = 1, then "CC" is a direct cause of "RR" or "RR" is a direct effect of "CC". If two variables directly affect a third variable, the first two variables are common causes, and the third variable is a collider. Evaluate the hypothesis based on the given partially presented as an adjacency matrix with the given Instruction.
A.4 Prompts for few-shot learning of $\text{C}^{2}\text{P}$
An example of given prompts for the few-shot learning process of $\text{C}^{2}\text{P}$ is as follows: